It is a word problem, I hope you can help,thank you in advance
Henk starts a new job selling tv sets. He sells 11 sets in the first week, 14 sets in the next, 17 in the next and so on in an arithmetic sequences. In what week does Henk sell his 2000th tv set?

Found 2 solutions by Cromlix, ikleyn:
Answer by Cromlix(4381)

Hi there,
Using the formula
a + (n - 1) d
a is the first number of the sequence
d is the difference between nos of TVs
number of TVs = 11 + (n - 1)3
2000 = 11 + 3n - 3
2000 = 8 + 3n
1992 = 3n
n = 664
On his 664 week Henk will sell his 2000th TV.
Hope this helps :-)

Answer by ikleyn(52812)

.
It seems to me that the question is not about "in which week does Henk sell 2000 tv sets per week".

It is about "in which week Henk will sell his 2000-th tv set".

Then you need operate with the sum of the arithmetic progression and look in which week this sum will get 2000.